If you are using 3.00% (mass/mass) hydrogen peroxide solution and you determine that the mass of solution required to reach the equivalence point is 5.125 g, how many moles of hydrogen peroxide molecules are present?
Answer #1

According to the given data

each 100gms of solution has 3gms of H2O2

then 5.125gms of solution has----------------------? gms of H2O2

mass of H_{2}O_{2}=\frac{5.125X3}{100}=0.153gms

molar \; mass of H_{2}O_{2}=34gms

hence

no. of \: moles \: of \; H_{2}O_{2}=\frac{mass}{molarmass}= \frac{0.153}{34}=0.0045moles

\therefore no. of \: moles \: of \; H_{2}O_{2}=0.0045moles